Posts

NPP Projection Report 9/6 Open Alec & Will

To do this evening/ was completed: Presentation Issues: Wyatt radioed about a dowser glitch at the start of Son of the White Mare , there was no picture on screen when the feature started. This happens occasionally for lens/masking changes--the dowser open cue goes through but the dowser does not physically respond to the cue, Will caught it and hit the dowser close/dowser open buttons on the projector which solved it. Guests asked us to rewind the film which I did. Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: The TEAL event(s) all went well! No serious issues; there was a bit of back and forth with the slides in T3, they ultimately wanted the screen black for the conversation and ceremony so the slides wouldn't bleed onto people's faces. Something we can improve next time. I also had the lights in T7 going down at 12:25 for the previews, as we discussed on Thursday, but I did not accurately reflect this on the cue sheet for today so there was a brief mom...

NPP Projection Close 9/5

To do this evening/ was completed: Built all A1 9/12 playlists with the exception of Demon Slayer-- the feature has not arrived yet. Built a playlist for the Animated Shorts Showcase but could not locate the intro/Q&A slide we need for it. Emailed Desmond. Inspected Longest Nite. It's in good condition. On house reels. Some of the film slightly exceeds the circumference of the house reels. Think it will be okay but just noting here. Also: I couldn't discern what type of soundtrack the film has. Nothing on the leader. It's either A-type or SR. IMDB says mono but the waveforms show otherwise.  Ingested Sisterhood of Traveling Pants to LMS. TEAL letters are up on the marquee. Conjuring sounds good at 5.5 with a full house (T1). Presentation Issues: Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Outgoing/Shipping (content): Print/video Inspection: Asset tracking doc updated/check...

NPP Projection Report 9/5 Admin

To do this evening/ was completed: The brunch marquee needs to be updated for the TEAL rental tonight.  Presentation Issues: Preshow Clips/Playlist: Built/scheduled both Jackass playlists. Trailer Updates: Special Events/ AV set up: Ordering/Receiving (posters & trailers): Ordering/Receiving (content): We received The Longest Nite print, this month's Sundays on Fire selection. Brought it to booth 1. Ingested Tropic Thunder . Outgoing/Shipping (content): Fantastic Planet was picked up. Print/video Inspection: Teched Jackass with Desmond, all set and the volume is noted. The print looks great. We also looked at almost all of the shorts for the NSF animation showcase, the only exception being Nuevo Rico. Seems some glitch happened in the DCP conversion, it came on screen for a second then went black/silent. Reconverting this one. We teched the content for the Thelma & Louise drag show. The curtain call DCP was just the song, no slide, so I remade that and ingested it. Las...

NPP Projection Close 9/4

To do this evening/ was completed: Presentation Issues: Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Put on US Open coverage on at Trees Lounge STB went smoothly and all gear used was broken down/returned to storage following the show Converted some mp3 files to DCP for the Thelma and Louise pre-show drag performance (Ramona Slick's audio file is missing from the shared drive) and transferred them to the LMS and theater 2 Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Outgoing/Shipping (content): Print/video Inspection: Finished inspecting Jackass: The Movie  and filled out its inspection report...the print is in mint condition and should look great on screen The Sisterhood of the Traveling Pants is continuing to convert to DCP in booth 1 and should finish sometime tomorrow evening Asset tracking doc updated/checked?: Maintenance/cleaning: Cue sheets updated/checked (every shift): Yes Marquee updated?: Organized letters for the TEA...

NPP Projection Report 9/4 Admin

To do this evening/ was completed: Presentation Issues: Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Spoons is all set up for tonight. Wireless mics are set on stands and tested. Yamaha mixer and spotlight controls are on the rollout table with the AJA extender. Spotlights and HDMI are working. I left everything on except the mics, Yamaha mixer, and spotlight control board. I don't know who broke down the PA system in T1 last so I'm noting this here: please remember to hang the XLR cable for the left speaker on the hook under the curtain, please don't leave it on the floor. The hook is way over to the left under the curtain, attached to the frame of the screen like the hook for the right speaker cable. Had a walkthrough with Devin from TEAL and Chenelle for the event on Saturday. We went through the run of show for each theater. I left a PA speaker set up in two of the three theaters they've rented so they're ready for Saturday. I'll be ru...

NPP Projection Close 9/3

To do this evening/ was completed: Brought Blonde Venus down to side entrance. Vicious Lips print has arrived, resting in T1. Updated 9/5 R playlists with OBAA  trailer #5. Fantastic Planet went smoothly. Print has been rewound on its shipping reels, ready for outgoing. No need for re-inspection report--as Alec mentioned in a recent email, only prints from the Academy Film Archive and LOC? need re-inspection. T7 last round--only two patrons, who inquired during preshow whether we could screen the non OCAP version of Honey Don't. We notified them that we could, there would just be a brief pause while the playlist loaded. Played fine. Presentation Issues: Preshow Clips/Playlist: Trailer Updates: Special Events/ AV set up: Ordering/Receiving (posters & trailers): Ordering/Receiving (content): Outgoing/Shipping (content): Print/video Inspection: Asset tracking doc updated/checked?: Maintenance/cleaning: Cue sheets updated/checked (every shift): Yes Marquee updated?: Posters updated...

Projection Admin 9/3

To do this evening/ was completed: Staff Screening info has been emailed out, scheduled on the LMS and blocked on the calendar. We've got additional staff screenings of Son of the White Mare per many requests.  Had our cinema meeting today. We're in good shape for the week we're about to enter, but we're still waiting on the performance pieces for Thelma and Louise from Malice/Ramona, but Desmond informed me they should be uploaded by today, EOD. If not, let's check in tomorrow.  We're just about ready for tech this Friday.  Presentation Issues: Preshow Clips/Playlist:  Trailer Updates: Please replace the 9/05 R rated playlists with One Battle After Another Trailer 5 (making sure conjuring isn't in there!) Special Events/ AV set up: Chenelle and I were to have a meeting with the TEAL folks today to hash out the last few bits of the private events this weekend, but it got pushed to tomorrow, so Alec will be on call for it. I sent him an email with my notes, ...